# Hot Rocks Survey III: A deep eclipse for LHS 1140c and a new Gaussian process method to account for correlated noise in individual pixels

### v1

@@ Appendix C Eclipse timing and eccentricity constraints
Given that the time of primary transit is already tightly constrained by transit observations, and our observations are the first to detect the eclipse of LHS 1140c, this should permit a tight constraint on the eccentricity parameter 666And in principle on other stellar parameters such as the stellar density (e.g. see Mahajan et al. 2024), although we do not explore this within this work. This is because the time gap from primary to secondary transit (ignoring light delay) is strongly dependent on . This time gap is given in Alonso (2018) as:
@@ C.1 Individual eclipse time constraints
We also tested fitting each eclipse completely independently using aperture extraction with a GP, as opposed to Section 4.3.1 where we used eclipse times from the other two eclipses to constrain each eclipse time. Figure 20 shows the constraints on the central eclipse times (given relative to a circular orbit) and eclipse depths for each of the three eclipses fit independently. We see that while each individual dataset does not strongly detect an eclipse, the highest probability central eclipse times all peak close to the expected time for a circular orbit and the eclipse depths are all closely matching. This consistency is strong evidence that our observed eclipses are real and not the result of systematics.
In addition to testing a fully independent fit to each of the three datasets, we also tested a joint-fit with a shared eclipse depth but separate eclipse times (using aperture extraction with GPs). This could be thought of as a joint-fit which assumes a circular orbit but accounts for transit timing variations (TTVs), although these are probably unlikely given the large separation between the two known planets in the system. Unfortunately, due to the low signal-to-noise of each individual eclipse, the central eclipse time is not strongly constrained by each individual dataset and so a uniform prior was chosen bounded within 30 minutes either side of the expected eclipse time for a circular orbit. We found that all three eclipses were favoured to occur a couple of minutes before the expected time with zero eccentricity, which explains the negative constraint when joint-fitting eccentricity parameters to all three datasets.
Table 10 gives the central eclipse time constraints for each eclipse from this joint-fit without adjusting for light delay. Figure 21 plots a histogram of the posterior for each of the central eclipse times. Note that Table 10 approximates these posteriors using the median value and upper and lower Gaussian uncertainties. The accuracy of these approximations is visualised by overlaying the probability distributions they produce in orange over the true posteriors in blue. The posteriors deviate substantially from symmetric Gaussian distributions but the different upper and lower uncertainties appear to do a reasonable job at tracing the true posterior.
Note that the constraints in Table 10 were used as priors when fitting individual eclipse depths as described in Section 4.3.1 and Appendix G. The way this was performed was by using the central eclipse time constraints from the two eclipses not being fit as priors on the expected eclipse time for the eclipse which was being fit (after adding/subtracting the correct number of periods between them). This helped constrain the individual eclipse fits which otherwise would have struggled to constrain the central eclipse time from a single dataset.
@@ Appendix D Tables of priors for both reduction analyses
The choice of priors chosen for various parameters in the primary and secondary reductions analyses are included in Tables 11 and 12.
Parameter | Distribution | Prior |
Uniform | [-, ] | |
Normal | ||
Normal | ||
Normal | ||
Normal | ||
Trunc. Normal | & | |
Normal | ||
Normal |
@@ Appendix E Atmospheres containing both SO2 and CO2
Our models suggest that the combination of SO2 at concentrations between 1 ppm - 1000 ppm in a CO2 background are ruled out to very low pressures of just 1 mbar to at least . Figure 22 summarises these constraints and highlights which compositions and pressures are ruled out at . Note that these models all include heat redistribution as modelled by the analytic parameterisation in Koll (2022). These results are intriguing as they suggest that eclipse photometry at 15m may be particularly sensitive to volcanic atmospheres. The addition of small concentrations of SO2 appears to make the CO2 absorption feature more prominent in the model spectra, likely due to the effect of SO2 on the temperature-pressure profile of the atmosphere.
@@ Appendix G Individual eclipse depths for different pixel-fits
While running pixel-fitting including the various components of pixel systematics as chosen in Section 3 and 4 such as including common systematics and independent pixel systematics, we also tested including more than just the 81 pixels which are 5px away from the centre of the PSF. We chose a larger radius of 8.5px to examine how this affects the results. Similar to optimal extraction, the method should be able to weight away from lower flux pixels so adding more pixels should always either tighten the constraints or leave them unchanged. We found that increasing from 5px away to 8.5px away (225 total pixels) had a negligible impact on the result other than slightly tightening the eclipse depth constraints. This was at a substantial computational cost as for each pixel we fit for the baseline flux and slope in flux , white noise amplitude , and height scale of independent pixel systematics (unless there is a shared height scale). Therefore this required fitting for more than 500 extra parameters and resulted in little change in the result. However, it is reassuring that our choice of pixels to include only had a small effect on the result. Note that for these pixel-fitting analyses a slightly larger convergence criterion of was used (instead of for the results presented in Section 4) in order to reduce computational costs.
@@ G.1 Testing different reduction steps
We tested varying the radius aperture between 4px and 6px instead of 5px but it had minimal effect on the eclipse depth as can be seen in Table 14. We also examined just running stage 1 and 2 of the pipeline with its default settings, using the jump step with the default threshold or skipping the jump step entirely. In addition, we tested skipping the emicorr step which corrects for electromagnetic interference in the raw data, skipping the flat-fielding step in stage 2 as well as including all groups (i.e. skipping the steps which exclude the first and last groups). We don’t find a dramatic difference for any of the various reductions we tested which suggests our results are not highly sensitive to these reduction choices.
@@ Appendix H Autocorrelation of pixel light curves
We can autocorrelate the pixel light curves for each eclipse to empirically inform our choice of kernel function. In practice this is somewhat complicated to implement as our pixel time-series can be thought of as a 3D data cube with two spatial dimensions and a time dimension (X, Y, t). Figure 24 attempts to examine correlations in this space. In the centre of the (5, 5) subplot is the autocorrelation of each pixel light curve with itself as a function of separation in time, averaged together across all pixels included within the pixel fit (5px from the PSF centre) and averaged across all three eclipse datasets. This is plotted in blue when performed on the real JWST data. To examine our kernel function choice we also took our best-fit pixel-fit to the real dataset (accounting for both common systematics and independent pixel systematics) and generated random draws of noise for each of the three real datasets using the covariance matrix generated by our kernel function. We could then autocorrelate this noise in the same manner as for the real data which was then plotted in orange. Finally, instead of performing this once for each eclipse and averaging together, the autocorrelation plotted in green was recovered by generating noise 100 times for each eclipse dataset and averaging together the autocorrelation, giving a better idea for how the kernel function converges for a lot of synthetic observations.
For the plot immediately left or right of the centre plot, the autocorrelation was taken between horizontally adjacent pixels as a function of separation in time. Similarly the pixels above and below the central plot are for vertically adjacent pixels and so on for diagonally adjacent pixels. The outer plots are all for pixels that are separated with one pixel gap between them i.e. two pixels apart. Note that for our kernel choice, we expect neighbouring pixels to show an anti-correlation in time as we expect them to be flux-conserved. This is clear in the green light curves are there has been one hundred covariance matrix draws to average over. For both the blue and orange light curves they only have three eclipse datasets to average over so they may only show moderate evidence of anti-correlation. However there does appear to be some sign of this anticorrelation, particularly in the adjacent neighbouring pixels where the blue light curves show signs of the V-shape seen in the green light curves. The pixels separated by more than one pixel show such weak correlation in all cases that there isn’t evidence of anything to adjust in our kernel function for pixels at this separation.

### v2

@@ Appendix C Eclipse timing and eccentricity constraints
Given that the time of primary transit is already tightly constrained by transit observations, and our observations are the first to detect the eclipse of LHS 1140c, this should permit a tight constraint on the eccentricity parameter 999And in principle on other stellar parameters such as the stellar density (e.g. see Mahajan et al. 2024), although we do not explore this within this work. This is because the time gap from primary to secondary transit (ignoring light delay) is strongly dependent on . This time gap is given in Alonso (2018) as:
@@ C.1 Individual eclipse time constraints
We also tested fitting each eclipse completely independently using aperture extraction with a GP, as opposed to Section 4.3.1 where we used eclipse times from the other two eclipses to constrain each eclipse time. Figure 20 shows the constraints on the central eclipse times (given relative to a circular orbit) and eclipse depths for each of the three eclipses fit independently. We see that while each individual dataset does not strongly detect an eclipse, the highest probability central eclipse times all peak close to the expected time for a circular orbit and the eclipse depths are all closely matching. This consistency offers strong evidence that our observed eclipses are real and not the result of systematics.
In addition to testing a fully independent fit to each of the three datasets, we also tested a joint-fit with a shared eclipse depth but separate eclipse times (using aperture extraction with GPs). This could be thought of as a joint-fit which assumes a circular orbit but accounts for transit timing variations (TTVs), although these are probably unlikely given the large separation between the two known planets in the system. Unfortunately, due to the low S/N of each individual eclipse, the central eclipse time is not strongly constrained by each individual dataset and so a uniform prior was chosen bounded within 30 minutes either side of the expected eclipse time for a circular orbit. We found that all three eclipses were favoured to occur a couple of minutes before the expected time with zero eccentricity, which explains the negative constraint when joint-fitting eccentricity parameters to all three datasets.
Table 10 gives the central eclipse time constraints for each eclipse from this joint-fit without adjusting for light delay. Figure 21 plots a histogram of the posterior for each of the central eclipse times. We note that Table 10 approximates these posteriors using the median value and upper and lower Gaussian uncertainties. The accuracy of these approximations is visualised by overlaying the probability distributions they produce in orange over the true posteriors in blue. The posteriors deviate substantially from symmetric Gaussian distributions but the different upper and lower uncertainties appear to do a reasonable job at tracing the true posterior.
We note that the constraints in Table 10 were used as priors when fitting individual eclipse depths as described in Section 4.3.1 and Appendix G. The way this was performed was by using the central eclipse time constraints from the two eclipses not being fit as priors on the expected eclipse time for the eclipse which was being fit (after adding/subtracting the correct number of periods between them). This helped constrain the individual eclipse fits which otherwise would have struggled to constrain the central eclipse time from a single dataset.
@@ Appendix D Tables of priors for both reduction analyses
The choice of priors chosen for various parameters in the primary and secondary reductions analyses are included in Tables 10 and 11.
@@ Appendix E Atmospheres containing both SO2 and CO2
Our models suggest that the combination of SO2 at concentrations between 1 ppm - 1000 ppm in a CO2 background are ruled out to very low pressures of just 1 mbar to at least . Figure 22 summarises these constraints and highlights which compositions and pressures are ruled out at . We note that these models all include heat redistribution as modelled by the analytic parameterisation in Koll (2022). These results are intriguing as they suggest that eclipse photometry at 15m may be particularly sensitive to volcanic atmospheres. The addition of small concentrations of SO2 appears to make the CO2 absorption feature more prominent in the model spectra, likely due to the effect of SO2 on the temperature-pressure profile of the atmosphere.
@@ Appendix G Individual eclipse depths for different pixel-fits
While running pixel-fitting including the various components of pixel systematics as chosen in Section 3 and 4 such as including common systematics and independent pixel systematics, we also tested including more than just the 81 pixels which are 5px away from the centre of the PSF. We chose a larger radius of 8.5px to examine how this affects the results. Similar to optimal extraction, the method should be able to weight away from lower flux pixels so adding more pixels should always either tighten the constraints or leave them unchanged. We found that increasing from 5px away to 8.5px away (225 total pixels) had a negligible impact on the result other than slightly tightening the eclipse depth constraints. This was at a substantial computational cost as for each pixel we fit for the baseline flux and slope in flux , white noise amplitude , and height scale of independent pixel systematics (unless there is a shared height scale). Therefore this required fitting for more than 500 extra parameters and resulted in little change in the result. However, it is reassuring that our choice of pixels to include only had a small effect on the result. We note that for these pixel-fitting analyses a slightly larger convergence criterion of was used (instead of for the results presented in Section 4) in order to reduce computational costs.
@@ G.1 Testing different reduction steps
We tested varying the radius aperture between 4px and 6px instead of 5px but it had minimal effect on the eclipse depth as can be seen in Table 14. We also examined just running stage 1 and 2 of the pipeline with its default settings, using the jump step with the default threshold or skipping the jump step entirely. In addition, we tested skipping the emicorr step which corrects for electromagnetic interference in the raw data, skipping the flat-fielding step in stage 2 as well as including all groups (i.e. skipping the steps which exclude the first and last groups). We did not find any dramatic difference for any of the various reductions we tested which suggests our results are not highly sensitive to these reduction choices.
@@ Appendix H Autocorrelation of pixel light curves
We can autocorrelate the pixel light curves for each eclipse to empirically inform our choice of kernel function. In practice this is somewhat complicated to implement as our pixel time series can be thought of as a 3D data cube with two spatial dimensions and a time dimension (X, Y, t). Figure 24 attempts to examine correlations in this space. In the centre of the (5, 5) subplot is the autocorrelation of each pixel light curve with itself as a function of separation in time, averaged together across all pixels included within the pixel fit (5px from the PSF centre) and averaged across all three eclipse datasets. This is plotted in blue when performed on the real JWST data. To examine our kernel function choice we also took our best-fit model from pixel-fitting the real dataset (accounting for both common systematics and independent pixel systematics) and generated random draws of noise for each of the three real datasets using the covariance matrix generated by our kernel function. We could then autocorrelate this noise in the same manner as for the real data which was then plotted in orange. Finally, instead of performing this once for each eclipse and averaging together, the autocorrelation plotted in green was recovered by generating noise 100 times for each eclipse dataset and averaging together the autocorrelation, giving a better idea for how the kernel function converges for a lot of synthetic observations.
For the plot immediately left or right of the centre plot, the autocorrelation was taken between horizontally adjacent pixels as a function of separation in time. Similarly the pixels above and below the central plot are for vertically adjacent pixels and so on for diagonally adjacent pixels. The outer plots are all for pixels that are separated with one pixel gap between them i.e. two pixels apart. We note that for our kernel choice, we expect neighbouring pixels to show an anti-correlation in time as we expect them to be flux-conserved. This is clear in the green light curves are there has been one hundred covariance matrix draws to average over. For both the blue and orange light curves they only have three eclipse datasets to average over so they may only show moderate evidence of an anti-correlation. However there does appear to be some sign of this anticorrelation, particularly in the adjacent neighbouring pixels where the blue light curves show signs of the V-shape seen in the green light curves. The pixels separated by more than one pixel show such a weak correlation in all cases that there is no evidence of anything to adjust in our kernel function for pixels at this separation.